This is a critical thinking exercise, that may or may not require algebra.
A Wagon Wheel is placed against a wall.One point on the edge of the wheel is 5 inches from the ground and 10 inches from the wall. What is the radius of the wheel?

Answer by nerdybill(7384)

This is a critical thinking exercise, that may or may not require algebra.
A Wagon Wheel is placed against a wall.One point on the edge of the wheel is 5 inches from the ground and 10 inches from the wall. What is the radius of the wheel?
.
Since the wheel is leaning against a wall, the "wall","ground", and the "wheel" forms a right triangle -- where the "diameter of the wheel" is the hypotenuse of the triangle.
.
Using Pythagorean theorem:
Let d = diameter of wheel
5^2 + 10^2 = d^2
25 + 100 = d^2
25 + 100 = d^2
125 = d^2
sqrt(125) = d
.
radius, of course, is half the diameter:
radius = sqrt(125)/2 = 5.59 inches